OpenStudy (anonymous):

Matt looks at the architectural plan of a four-walled room in which the walls meet each other at right angles. The length of one wall in the plan is 19 inches. The length of the diagonal of the floor of the room in the plan is approximately 20.62 inches. Is the room in the shape of a square? Explain how you determined your answer. Show all your work.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

if the shape of the room is a square, then I can write this formula: \(\huge d=\sqrt 2 \cdot l\) where \(\Large l\) is the length of the side of the room, and \(\Large d\) is the length of the diagonal of the room. Now, what is: \(\huge \sqrt2 \cdot 19=...?\)

OpenStudy (anonymous):

6.1 @Michele_Laino

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

hint: we have: \[\huge \sqrt 2 \times 19 \cong 1.414 \times 19 = ...?\]

OpenStudy (michele_laino):

@Marc1313

OpenStudy (anonymous):

26.866 @Michele_Laino